OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates ordered pair 7 comma 2 pi divided by 3.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

\[\large (7, \frac{ 2\pi }{ 3 })\]

jimthompson5910 (jim_thompson5910):

\[\large (r,\theta) = (7, \frac{ 2\pi }{ 3 })\] So, \[\large r = 7\] \[\large \theta = \frac{ 2\pi }{ 3 }\]

jimthompson5910 (jim_thompson5910):

they want you to convert to (x,y) form use these formulas \[\large x = r*\cos(\theta)\] \[\large y = r*\sin(\theta)\]

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np